# Researcher Population Pyramids for Tracking Global Demographic and Gender Trajectories

### v1

@@ Abstract
The sustainability of the global academic ecosystem relies on researcher demographics and gender balance, yet assessing these dynamics in a timely manner for policy is challenging. Here, we propose a researcher population pyramids framework for tracking global demographic and gender trajectories using publication data. This framework provides a timely snapshot of historical and present demographics and gender balance, revealing three contrasting research systems: Emerging systems (e.g., Arab countries) exhibit high researcher inflows with widening gender gaps in cumulative productivity; Mature systems (e.g., the United States) show modest inflows with narrowing gender gaps; and Rigid systems (e.g., Japan) lag in both. Furthermore, by simulating future scenarios, the framework makes potential trajectories visible. If 2023 demographic patterns persist, Arab countries’ systems could resemble mature or even rigid ones by 2050. Our framework provides a robust diagnostic tool for policymakers worldwide to foster sustainable talent pipelines and gender equality in academia.
Keywords: science of science; sustainability of the academic ecosystem; research careers; gender equality; cross-national analysis
@@ 1 Introduction
The pursuit of global sustainability–addressing grand challenges such as climate change, public health crises, and equitable economic development–is inextricably linked to the vitality of the global research enterprise. Sustained innovation, driven by a robust and dynamic academic ecosystem, provides the essential knowledge, technologies, and solutions required to navigate these complex issues [1]. However, the role of this ecosystem extends beyond immediate scientific discoveries; it is responsible for the long-term reproduction of research talent and the continuity of knowledge itself [2]. A disruption in this human capital pipeline, encompassing both the flow of researchers from early-career training to senior-level expertise and the diversity within that flow, undermines a society’s capacity to address not only current problems but also future, unforeseen challenges [3, 4, 5]. Ensuring the sustainability of the academic ecosystem is not merely an internal concern for the research community but a fundamental prerequisite for the long-term well-being and resilience of global society as a whole [6].
@@
The sustainability of the academic ecosystem, in turn, is fundamentally determined by the demographic composition of its researcher population [7, 8, 9]. This composition across career stages–early-career, mid-career, and senior-career researchers–directly influences the quality and quantity of research outputs within countries. Doctoral students and postdoctoral researchers constitute a significant part of the pipeline for future researchers, and individuals’ productivity during postdoc periods correlates with subsequent career continuity [10]. Importantly, across all career stages, researchers exhibit diverse productivity patterns [11], reflected in considerable variation in the timing of their most impactful research, which can occur at any stage of a research career [12]. Mid-career and senior-career researchers also play critical mentoring roles, thus facilitating the effective transfer of scientific expertise to younger generations [13, 14]. In addition to ensuring the researcher population across career stages, achieving gender diversity in academia remains a persistent global challenge [15, 16, 17, 18, 19, 20], as gender-diverse research environments foster broader research questions, varied perspectives, and enhanced scientific creativity [21, 22]. Therefore, developing robust methods for the systematic analysis of global researcher demographics and gender balance has become a critical challenge for science policymakers and educational administrators worldwide.
@@
Reconstructing researchers’ publishing careers from bibliographic data has enabled broader international and longitudinal analyses of research systems and research careers, including productivity patterns [12, 11], collaboration patterns [23, 24], research impacts [25, 26], structural imbalances [15, 18], and many more [27, 28, 2]. Quantifying global researcher demographics across career stages, gender, and national contexts from bibliometric data, however, still poses a significant methodological challenge: reliably determining whether researchers’ publishing careers are ongoing in specific years. This challenge stems from the diversity in publication intervals among researchers and technical barriers in identifying researchers’ publications based on bibliographic data [18, 29, 30]. This uncertainty necessitates a substantial time lag when assessing researchers’ career characteristics, including total productivity and publishing-career length, which often rely on the definition of their publishing-career end [18]. These delays are particularly problematic for rapidly evolving research systems, where timely insights are essential for informing effective research and educational policies.
@@
Here, we propose a researcher population pyramids framework for mapping and comparing global researcher demographics across publishing careers, gender, and national contexts. Based on author and publication data from a large-scale bibliographic database, we deploy country- and gender-specific thresholds of inter-publication intervals to identify “active” authors (i.e., those whose publishing careers are ongoing) at a specific year. Within this framework, we define the cumulative productivity of an active author as the number of their most recent, uninterrupted publications up to a given year. Unlike the total productivity and publishing-career length, cumulative productivity can be measured without waiting for a publishing career to end. We construct population pyramids by counting active authors according to their cumulative productivity in a given year by country and gender. This framework enables systematic international comparisons of researcher demographics and gender balance across time. For future contexts, we simulate potential trajectories based on plausible transition scenarios derived from observed trends.
@@ 2.1 Trends in Researcher Populations and Gender Balance
We analyze 15,104,205 gender-assigned authors, derived from 151,905,632 publication records. These authors are affiliated with 58 countries for which Naive Bayes gender classifiers achieved high accuracy in two benchmarks (see Methods and Supplementary Sections S1 and S2). For comparison, we focus on two country groups: nine Arab countries (e.g., Egypt and Tunisia) and ten reference countries (e.g., Japan and the United States; see Fig. 1 for full lists). We selected these countries to ensure a sufficiently large sample of gender-assigned authors (e.g., at least 9,000 per gender per country) and to represent regional diversity relevant for global comparisons. The comparison is primarily motivated by their contrasting developmental trajectories. Some Arab countries, like Egypt and Tunisia, experienced significant socio-political transformations during and after the Arab Spring (2010–2012) [31, 32]. Given these systemic changes and their notable increase in research output [33, 34, 35, 36, 37], we hypothesize that Arab countries are undergoing rapid shifts in researcher demographics and gender balance. By contrast, reference countries, with their established research infrastructures and stable funding mechanisms [36], serve as demographic benchmarks for comparison.
Figure 1: Trends in researcher population and gender balance from 2000 to 2023. Panels (a) and (b): Annual number of unique authors who published at least one paper, affiliated with (a) Arab and (b) reference countries. Plotted on a logarithmic scale. Panels (c) and (d): Annual proportion of female authors affiliated with (a) Arab and (b) reference countries. The gray shaded area marks the period of the Arab Spring (2010–2012).
@@
Furthermore, to explore the impacts of specific socio-political events, we examined the trends in researcher populations and gender balance in Egypt and Tunisia, countries that underwent substantial political transitions during the Arab Spring. Our analysis, which specifically compared pre- and post-Arab Spring periods (excluding 2010–2013), revealed contrasting effects: Egypt saw an increased growth rate in female author proportions while its researcher population growth remained stable, whereas Tunisia experienced declines in both metrics (see Supplementary Section S3.3). These findings highlight how major events can influence researcher demographics and gender balance, partially consistent with previous studies on the impacts of the Arab Spring on regional academia [33, 39].
@@ 2.2 Cumulative Productivity and Population Pyramids
Figure 2: Researcher population pyramids in 2010, 2023, and 2050 for Egypt, Tunisia, Japan, and the United States. Cumulative productivity is plotted on the vertical axis, while the number of active authors for each gender is displayed on a logarithmic horizontal axis. If the number of authors with cumulative productivity $k$ is zero, the bar for $k$ is not displayed.
To provide a more detailed understanding of researcher demographics, gender balance, publishing careers, and their evolution, we introduce a researcher population pyramids framework. This framework allows for the multi-dimensional visualization of researcher populations across publishing-career stages and genders. We construct these pyramids by counting active authors (i.e., those whose publishing careers are ongoing) in a given year by their cumulative productivity (i.e., the number of their most recent, uninterrupted publications up to that year; see Methods). Unlike existing metrics that rely on a definitive career end, cumulative productivity can be measured without waiting for a publishing career to end. We also simulate a plausible future demographic pattern by extending the 2023 pyramid forward (see Methods); our projections should be interpreted as scenarios based on 2023 trends.
@@
In 2010, both Egypt and Tunisia exhibited relatively limited numbers of authors across publishing-career stages. Between 2010 and 2023, both countries experienced substantial expansion in their researcher populations, driven primarily by marked increases in early- and mid-career authors. Looking towards 2050, our projections indicate substantial growth for Egypt, with both the number of authors (horizontal expansion) and cumulative productivity levels (vertical growth) continuing to expand. This suggests Egypt exemplifies a research system undergoing rapid demographic expansion. In contrast, Tunisia’s 2050 projection suggests a more moderate overall growth characterized by a slowing inflow of early-career authors and an increasingly stable demographic profile. This pattern indicates Tunisia’s research system is transitioning toward a more mature phase than Egypt’s.
@@
The United States’ 2010 population pyramid showed a well-balanced distribution of researchers across all publishing-career stages, despite a moderate underrepresentation of women. This stable structure was maintained through 2023, with observable improvements in gender imbalance. Projections for 2050 suggest continued structural stability and only modest expansion. This trajectory–reflecting a research system that may be reaching saturation but appears demographically sustainable–presents a stark contrast to the dynamic growth observed in Egypt. Consequently, the United States exemplifies a mature, steady-state research system.
@@
Underlying the divergent shapes of the population pyramids, we found that countries vary greatly in their researcher inflow–the proportion of newly active authors among all active authors affiliated with a country in a given year. For 2023, Arab countries showed substantially higher researcher inflow than reference countries (Fig. 3). Furthermore, the share of women among newly active authors achieves near gender parity (defined here as 45–55% female representation) in most Arab countries and many reference countries. However, notable exceptions exist: Lebanon, Tunisia, and the United Arab Emirates exhibited slightly higher female shares than the defined parity range. In contrast, Japan consistently exhibited the lowest proportion of female newly active authors among all countries considered.
@@
To contextualize the observed inflow dynamics and future projections for diverse research systems, we examine the long-term historical evolution of research populations in Japan and the United States from 1970 to 2000 (Fig. 4). In 1970, Japan’s pyramid featured a narrow base with very few senior female authors. By 1980, Japan’s pyramid structure closely resembled that of the United States in 1970, suggesting Japan lagged about a decade behind the United States in its demographic development. However, the United States pyramid expanded rapidly thereafter; by the 1990s and 2000s, its shape mirrored that of Japan in 2023 (see also Fig. 2). This comparison implies that Japan now lags more than 20 years behind the United States in Japan’s population structure, indicating the development gap has widened. These retrospective pyramids provide a valuable reference point for interpreting the potential future trajectories of today’s rapidly growing research systems, such as those in the Arab countries.
@@
The preceding results lead us to position the 58 countries along two key dimensions: researcher inflow and gender gap in cumulative productivity (Fig. 5). First, all of these countries exhibit a negative gender gap in cumulative productivity (i.e., the mean for women is lower than for men), which is largely consistent with previous results on gender gaps in total productivity and publishing-career length across countries [18]. Second, the trajectories from 2010 to 2023 (gray arrows) reveal three distinct patterns of demographic evolution: Emerging systems, exemplified by most Arab countries, are characterized by high researcher inflow and a widening gender gap in cumulative productivity (arrows moving to the left). Mature systems, such as the United States, exhibit lower researcher inflow but a narrowing gender gap (arrows moving down and to the right). Rigid systems, typified by Japan, are marked by both the lowest researcher inflow and the largest, most persistent gender gap (minimal arrow movement to the right).
@@ 3 Discussion
We examined global researcher demographics through the lens of population pyramids, identifying three distinct types of research systems: (i) Emerging systems, exemplified by current Arab countries, which exhibit large inflows of newly active authors alongside widening gender gaps in cumulative productivity; (ii) Mature systems, such as the United States, characterized by moderate inflows and gradually narrowing gender gaps; and (iii) Rigid systems, such as Japan, which lag in both researcher inflows and progress on gender equality.
@@
Our approach has additional limitations. First, the cumulative productivity metric has inherent limitations. It does not capture non-publishing contributions (e.g., teaching, mentorship, and administrative roles), nor does it fully account for career breaks such as parental leave or work in other sectors. Furthermore, as cumulative productivity is based on the number of publications, our analysis may inadvertently undervalue researchers in disciplines with a lower annual publication volume. Second, our analysis excludes several major research-producing countries, including China, India, and South Korea, due to the current limitations of name-based gender inference for these populations. Future work should develop improved gender inference methods tailored to names in these countries and conduct dedicated demographic studies of their research systems. Such efforts are essential for a more inclusive global understanding of researcher population dynamics.

### v2

# Researcher Population Pyramids: Tracking Demographic and Gender Trajectories Across Countries

@@ Abstract
The sustainability of the academic ecosystem relies on researcher demographics and gender balance, yet assessing these dynamics in a timely manner for policy is challenging. Here, we propose a researcher population pyramid framework for tracking demographic and gender trajectories across countries using publication data. We provide a timely snapshot of historical and present demographics and gender balance across 58 countries, revealing three contrasting patterns among research systems: Emerging systems (e.g., Arab countries) exhibit high researcher inflows with widening gender gaps in cumulative productivity; Mature systems (e.g., the United States) show modest inflows with narrowing gender gaps; and Rigid systems (e.g., Japan) lag in both. Furthermore, by simulating future scenarios, the framework makes potential trajectories visible. If 2023 demographic patterns persist, Arab countries’ systems could resemble mature or even rigid ones by 2050. Our framework provides a robust diagnostic tool for policymakers worldwide to foster sustainable talent pipelines and gender equality in academia.
Significance Statement
We introduce a researcher population pyramid framework for tracking cross-national demographics and gender balance. Applying this framework to 58 countries, we uncover three contrasting patterns in academic systems—which we term “Emerging”, “Mature”, and “Rigid”—each facing unique challenges in researcher inflow and gender equity. For example, while many Arab countries show rapid growth in researcher inflow, they also exhibit widening gender gaps in cumulative productivity, signaling a critical policy window to foster equitable academic futures. This framework provides policymakers worldwide with a data-driven diagnostic tool to promote long-term sustainability and gender equality in academia.
Keywords: science of science; academic sustainability; research ecosystem; research careers; gender equality
@@ 1 Introduction
The pursuit of global sustainability—addressing grand challenges such as climate change, public health crises, and equitable economic development—is inextricably linked to the vitality of the global research enterprise. Sustained innovation, driven by a robust and dynamic academic ecosystem, provides the essential knowledge, technologies, and solutions required to navigate these complex issues [1]. However, the role of this ecosystem extends beyond immediate scientific discoveries; it is responsible for the long-term reproduction of research talent and the continuity of knowledge itself [2]. A disruption in this human capital pipeline, encompassing both the flow of researchers from early-career training to senior-level expertise and the diversity within that flow, undermines a society’s capacity to address not only current problems but also future, unforeseen challenges [3, 4, 5]. Ensuring the sustainability of the academic ecosystem is not merely an internal concern for the research community but a fundamental prerequisite for the long-term well-being and resilience of global society as a whole [6].
@@
The sustainability of the academic ecosystem, in turn, is fundamentally determined by the demographic composition of its researcher population [7, 8, 9]. This composition across career stages—early-career, mid-career, and senior-career researchers—directly influences the quality and quantity of research outputs within countries. Doctoral students and postdoctoral researchers constitute a significant part of the pipeline for future researchers, and individuals’ productivity during postdoctoral periods correlates with subsequent career continuity [10]. Importantly, across all career stages, researchers exhibit diverse productivity patterns [11], reflected in considerable variation in the timing of their most impactful research, which can occur at any stage of a research career [12]. Mid-career and senior-career researchers also play critical mentoring roles, thus facilitating the effective transfer of scientific expertise to younger generations [13, 14]. In addition to ensuring a balanced researcher population across career stages, achieving gender diversity in academia remains a persistent global challenge [15, 16, 17, 18, 19, 20, 21, 22, 23], as gender-diverse research environments foster broader research questions, varied perspectives, and enhanced scientific creativity [24, 25]. Therefore, developing robust methods for the systematic analysis of researcher demographics and gender balance across countries has become a critical challenge for science policymakers and educational administrators worldwide.
@@
Reconstructing researchers’ publishing careers from bibliographic data has enabled broader international and longitudinal analyses of research systems and research careers, including productivity patterns [12, 11], collaboration patterns [26, 27, 28, 29], research impacts [30, 31], structural imbalances [15, 19], and many more [32, 33, 2]. Quantifying researcher demographics across career stages, gender, and national contexts from bibliometric data, however, still poses a significant methodological challenge: reliably determining whether researchers’ publishing careers are ongoing in specific years. This challenge stems from the diversity in publication intervals among researchers and technical barriers in identifying researchers’ publications based on bibliographic data [19, 34, 35]. This uncertainty, as assessments of total productivity and publishing-career length often rely on defining a career end [19], necessitates a substantial time lag when evaluating researchers’ career characteristics. These delays are particularly problematic for rapidly evolving research systems, where timely insights are essential for informing effective research and educational policies.
@@
Here, we propose a researcher population pyramid framework for mapping and comparing researcher demographics across publishing careers, gender, and national contexts. Based on author and publication data from a large-scale bibliographic database, we deploy country- and gender-specific thresholds of inter-publication intervals to identify “active” authors (i.e., those whose publishing careers are ongoing) in a specific year. Within this framework, we define the cumulative productivity of an active author as the number of their most recent, uninterrupted publications up to a given year. Unlike the total productivity and publishing-career length, cumulative productivity can be measured without waiting for a publishing career to end. We construct population pyramids by counting active authors according to their cumulative productivity in a given year by country and gender. This framework enables systematic international comparisons of researcher demographics and gender balance across time. For future contexts, we simulate potential trajectories based on plausible transition scenarios derived from observed trends. Applying this diagnostic tool to 58 countries, we demonstrate its utility by revealing three contrasting evolutionary patterns—Emerging, Mature, and Rigid—exemplified by Arab countries, the United States, and Japan, respectively.
@@ 2 Results
Figure 1: Trends in researcher population and gender balance from 2000 to 2023. Panels (a) and (b): Annual number of unique authors who published at least one paper, affiliated with (a) Arab and (b) reference countries. Plotted on a logarithmic scale. Panels (c) and (d): Annual proportion of female authors affiliated with (a) Arab and (b) reference countries.
@@ 2.1 Trends in Researcher Populations and Gender Balance
We analyze 14,745,796 gender-assigned authors, derived from 151,905,632 publication records. These authors are affiliated with 58 countries for which Naive Bayes gender classifiers achieved high accuracy in two benchmarks (see Methods and Supplementary Sections S1 and S2). For comparison, we focus on two country groups. The nine Arab countries are Algeria, Egypt, Iraq, Jordan, Lebanon, Morocco, Saudi Arabia, Tunisia, and the United Arab Emirates. The ten reference countries are Australia, Canada, France, Germany, Italy, Japan, Spain, Sweden, the United Kingdom, and the United States. We selected these countries to ensure a sufficiently large sample of gender-assigned authors (e.g., at least 9,000 per gender per country) and to represent regional diversity relevant for cross-national comparisons. The comparison is primarily motivated by their contrasting developmental trajectories. In particular, Arab countries experienced significant socio-political transformations during and after the Arab Spring (2010–2012) [36, 37]. Given these systemic changes and their notable increase in research output [38, 39, 40, 41, 42], we hypothesize that Arab countries are undergoing rapid shifts in researcher demographics and gender balance. By contrast, reference countries, with their established research infrastructures and stable funding mechanisms [41], serve as demographic benchmarks for comparison.
@@ 2.2 Cumulative Productivity and Population Pyramids
Figure 2: Researcher population pyramids in 2010, 2023, and 2050 for Egypt, Tunisia, Japan, and the United States. The number of active authors for each gender is displayed on a logarithmic horizontal axis. Total active author counts (NN) are provided for each panel. The 2050 pyramids and their corresponding counts ($\approx$) are projections based on 2023 trends.
To provide a more detailed understanding of researcher demographics, gender balance, publishing careers, and their evolution, we introduce a researcher population pyramid framework. This framework allows for the multi-dimensional visualization of researcher populations across publishing-career stages and genders. We construct these pyramids by counting active authors (i.e., those whose publishing careers are ongoing) in a given year by their cumulative productivity (i.e., the number of their most recent, uninterrupted publications up to that year; see Methods). Unlike existing metrics that rely on a definitive career end, cumulative productivity can be measured without waiting for a publishing career to end. We also simulate a plausible future demographic pattern by extending the 2023 pyramid forward (see Methods); our projections should be interpreted as scenarios based on 2023 trends.
@@
In 2010, both Egypt and Tunisia exhibited relatively limited numbers of authors across publishing-career stages. Between 2010 and 2023, both countries experienced substantial expansion in their researcher populations, driven primarily by marked increases in early- and mid-career authors. Looking towards 2050, our projection scenario suggests substantial growth for Egypt, with both the number of authors (horizontal expansion) and cumulative productivity levels (vertical growth) continuing to expand. This suggests Egypt exemplifies a research system undergoing rapid demographic expansion. In contrast, Tunisia’s 2050 projection suggests a more moderate overall growth characterized by a slowing inflow of early-career authors and an increasingly stable demographic profile. This pattern indicates Tunisia’s research system is transitioning toward a more mature phase than Egypt’s.
@@
The United States’ 2010 population pyramid showed a well-balanced distribution of researchers across all publishing-career stages, despite a moderate underrepresentation of women. This stable structure was maintained through 2023, with observable improvements in gender imbalance. Projections for 2050 suggest continued structural stability and only modest expansion. This trajectory—reflecting a research system that may be reaching saturation but appears demographically sustainable—presents a stark contrast to the dynamic growth observed in Egypt. Consequently, the United States exemplifies a mature, steady-state research system.
Figure 4: Researcher population pyramids for the aggregate of the nine Arab countries in 2000, 2005, 2010, 2015, 2020, and 2023.
@@
Underlying the divergent shapes of the population pyramids, we found that countries vary greatly in their researcher inflow—defined as the proportion of newly active authors among all active authors affiliated with a country in a given year. For 2023, Arab countries showed substantially higher researcher inflow than reference countries (Fig. 3). Furthermore, the share of women among newly active authors achieves near gender parity (defined here as 45–55% female representation) in most Arab countries and many reference countries. However, notable exceptions exist: Lebanon, Tunisia, and the United Arab Emirates exhibited slightly higher female shares than the defined parity range. In contrast, Japan consistently exhibited the lowest proportion of female newly active authors among all countries considered.
To provide a longitudinal perspective on the structural evolution associated with this researcher inflow, we examined the evolution of the aggregate researcher population pyramid, constructed by pooling all active authors affiliated with any of the nine Arab countries, from 2000 to 2023 (Fig. 4). In 2000, the pyramid exhibited a limited number of authors across all cumulative productivity levels. However, over the subsequent two decades, these countries experienced a marked horizontal widening of the early-career base, reflecting the substantial inflow of researchers. By 2023, this sustained inflow had begun to translate into vertical growth, with an increasing number of authors reaching higher levels of cumulative productivity. This structural shift demonstrates how our framework can visualize the transition of the “Emerging” research system.
@@
To contextualize the observed inflow dynamics and future projections for diverse research systems, we examine the long-term historical evolution of research populations in Japan and the United States from 1970 to 2000 (Fig. 5). In 1970, Japan’s pyramid featured a narrow base with very few senior female authors. By 1980, Japan’s pyramid structure closely resembled that of the United States in 1970, suggesting Japan lagged about a decade behind the United States in its demographic development. However, the United States pyramid expanded rapidly thereafter; by the 1990s and 2000s, its shape mirrored that of Japan in 2023 (see also Fig. 2). This comparison implies that Japan now lags more than 20 years behind the United States in its population structure, indicating the development gap has widened. These retrospective pyramids provide a valuable reference point for interpreting the potential future trajectories of today’s rapidly growing research systems, such as those in the Arab countries.
@@
The preceding results lead us to position the 58 countries along two key dimensions: researcher inflow and gender gap in cumulative productivity (Fig. 6). First, all of these countries exhibit a negative gender gap in cumulative productivity (i.e., the mean for women is lower than for men), which is largely consistent with previous results on gender gaps in total productivity and publishing-career length across countries [19]. Second, the trajectories from 2010 to 2023 (gray arrows) suggest three broad, qualitative patterns of demographic evolution. For descriptive purposes, we heuristically label these patterns as follows:
  • “Emerging”: Characterized by high researcher inflow and a widening gender gap in cumulative productivity (e.g., most Arab countries; arrows moving to the left).
  • “Mature”: Characterized by lower researcher inflow but a narrowing gender gap (e.g., the United States; arrows moving down and to the right).
  • “Rigid”: Characterized by low researcher inflow and a large, persistent gender gap (e.g., Japan; minimal arrow movement).
It is important to note that these labels are intended as a descriptive heuristic to facilitate interpretation. This qualitative classification is based on the observed positions and trajectories of countries within the specific two-dimensional space defined by the two metrics. We also note that while most of the Arab countries analyzed exhibited a decrease in their researcher inflow from 2010 to 2023 (indicated by the downward direction of the gray arrows in Fig. 6), they consistently maintained researcher inflow above the global median (the horizontal dashed line). This trend suggests that while the initial rapid growth of their researcher populations has begun to stabilize as the total number of active authors increases, these systems continue to exhibit a substantially higher proportion of newly active authors relative to mature or rigid systems.
@@ 3 Discussion
We examined researcher demographics across 58 countries through the lens of population pyramids, observing three broad patterns among research systems. We heuristically label these as follows: (i) “Emerging”, exemplified by current Arab countries, which exhibit large inflows of newly active authors alongside widening gender gaps in cumulative productivity; (ii) “Mature”, such as the United States, characterized by moderate inflows and gradually narrowing gender gaps; and (iii) “Rigid”, such as Japan, which lag in both researcher inflows and progress on gender equality.
@@
Regarding the validity of using cumulative productivity as a proxy for career stages, we note that this metric is a functional proxy based on researchers’ uninterrupted publication sequences rather than total productivity or publishing-career length. While total productivity and publishing-career length serve as measures of academic careers [19, 34, 35], they often require identifying a definitive career end and thus entail a substantial time lag when evaluating researchers’ career characteristics [19]. We employed cumulative productivity in our framework because it provides a timely snapshot of the active publishing workforce, serving as a data-driven diagnostic tool for monitoring research ecosystems.
While we heuristically labeled the bins of cumulative productivity as “early-career” (1–10 cumulative publications), “mid-career” (11–50), and “senior-career” (51 or more) to facilitate international comparisons, this classification may introduce potential biases in assessing researcher demographics and gender balance. First, despite the heterogeneity of publishing-career trajectories [11], cumulative productivity may conflate current publishing-activity levels with “academic age” (e.g., years since the first publication [59]). This, for instance, might categorize publishing-active researchers with a short academic age (e.g., those with 30 publications within five years of their first publication) as “mid-career”, potentially overestimating their academic careers. Second, publication practices vary significantly across disciplines. For example, average total productivity in a research career in biology (Male: 16.56, Female: 10.31) was substantially higher than in mathematics (Male: 7.13, Female: 5.55) in a previous analysis [19]. Our pyramids might reflect a complex interplay between national disciplinary composition and the underlying demographic distribution of publishing momentum. Nevertheless, our additional analysis across four major research domains reveals that the qualitative characteristics of the Emerging, Mature, and Rigid research systems remain largely consistent within Egypt, Tunisia, Japan, and the United States (see Supplementary Section S8 for details). Third, the cumulative productivity metric is sensitive to career interruptions; a career break (e.g., due to parenthood [51]) that exceeds the gender- and country-specific threshold resets the count to one in our framework. Although we defined these thresholds to accommodate diverse publication paces across countries and genders, the metric still prioritizes recent and continuous publishing momentum, potentially misclassifying experienced researchers with non-linear career trajectories as “early-career”.
We acknowledge that these methodological choices and their potential biases might shape the interpretation of our results. First, the observed gender gap in the “senior-career” stage may reflect a disparity in publishing continuity. In “Emerging” systems, widening gaps in cumulative productivity may signify a “publishing-continuity gap”, where female researchers—who disproportionately face career interruptions [51]—could be assigned to lower cumulative-productivity bins. Consequently, our findings should be interpreted as highlighting the structural challenges women face in maintaining uninterrupted publication momentum. Second, the classification of research systems into “Emerging”, “Mature”, and “Rigid” systems reflects the aggregate dynamics of active publication momentum. For instance, a “Rigid” system like Japan’s likely reflects not only low researcher inflow but also structural barriers or disciplinary compositions that prevent mid- and senior-career researchers from sustaining high publication momentum.
Our approach has additional limitations. First, the geographical scope of our analysis is constrained by the performance of name-based gender inference. We had to exclude several major research-producing countries, including China, India, and South Korea, where the classifier’s accuracy was insufficient. This insufficient accuracy result is partially consistent with recent work demonstrating that the error rates of name-based gender inference are not uniformly distributed [60]. Second, even among the countries included in our analysis, the gender assignment rate (i.e., the proportion of unique authors for whom a binary gender were inferred with high confidence) varies considerably (see Supplementary Table S7). For some countries, the rate is below 50%; notably, for Iraq, a key country in our analysis of the Arab countries, the rate is 45.8%. While we retained countries with lower rates in our analysis to maintain broad geographical coverage, their specific results should be interpreted carefully, as the subset of gender-assigned authors may not be representative of the entire researcher population in those countries. Third, while we focus on the longitudinal visualization and classification of research systems in this study, our framework may be useful for quantitatively investigating the causal impact of specific sociopolitical events or policy interventions on researcher demographics. Future work could use this diagnostic tool to assess whether historical events (e.g., the Arab Spring in Arab countries [38, 44]) reshaped the generational profiles of academic systems or whether these systems remained resilient to such events. Finally, our projections to 2050 are based on a strong stationarity assumption that the inflow in 2023 and transition patterns observed from 2022 to 2023 will remain constant in the future. More sophisticated forecasting models that incorporate uncertainty, scenario variations (e.g., changes in researcher inflow), and structural breaks warrant future work.